# Creative and geometric times in physics, mathematics, logic, and philosophy

### v1

@@ V Relation to the philosophy of time
Before discussing in detail how our proposal of creative and geometric times relates to some positions in the philosophy of time, we would like to stress that, as most physicists, we are not taking a Humean position. On the contrary, we believe that laws of nature are real causal connections (they have an ontological status) and that they guide the evolution of things of the world at different times. Note that these laws may not connects states at different times by necessity (determinism), but only state a propensity [4]. In either case, this puts an emphasis on the role of time in physical laws.
@@
In the vast literature devoted to the philosophy of time (see, e.g., [31, 32] and references therein), one finds a main separation into two camps, following the seminal works of J. M. E. McTaggart [33]. The two camps are customarily referred to as A-theory and B-theory of time. While we acknowledge that the distinction in A- and B-theories presents a variety of modification and it may not even encapsulate the most striking features of time in physics (e.g., directionality, openness of the future, objective becoming, etc.), we cannot avoid drawing some parallel with our distinction between creative and geometric time, respectively.
In fact, starting in reversed order, B-theorists regard positioning in time in dyadic relational terms: given two events $E_{1}$ and $E_{2}$, it is either the case that $E_{1}$ is before $E_{2}$, or simultaneous to $E_{2}$, or after $E_{2}$ [34]. Clearly, these relations are somehow ``````` `static”, i.e., they are not tensed and hold unchanged since ever and forever. For example, according to the B-theory, the extinction of the dinosaurs has always been, is and will always remain before the French revolution. Note however, that this relation seems to us to necessarily presume determinism, otherwise, in an indeterministic world, it would have not been true at a time earlier than the dinosaurs’ extinction itself actualized, because that extinction was not necessary in the first place (nor was the French revolution). At the remote past time in which both events ``````` `dinosaurs’ extinction” and ``````` `French revolution” were only potential events, their relation of before-after was also indeterminate. After the actualization of dinosaurs’ extinction but without an actualization of the French revolution, one rules out the possibility that the French revolution is before (or simultaneous to) the dinosaurs’ extinction, but it remains in general indeterminate whether the French revolution will ever actualize (although this may have a determined propensity [4]). Hence, B-theorists – who take the before-after relation as fixed – don’t derive determinism but assume it from the outset.
Regarding the metaphysical status of time, the B-theory naturally fits to the spacetime of (deterministic) classical theory and the theory of relativity, where events are organised on a 4-dimensional manifold. Moreover, B-theorists uphold Eternalism – i.e., the doctrine that all events are temporally on an equal footing and their relation of being in the past, present or future of each other – merely reflects their placement on the manifold [35]. Clearly, the concept of geometric time as expressed above falls into this B-theory camp.
@@
On the contrary, A-theorists uphold the view that characterizes the positioning in time with a monadic attribution: given an event $E$, it is either the case that $E$ is present, or past, or future [34]. For A-theorists change really happens: ``````` `there is a way reality is (now, presently) which is complete [i.e., maximal given what exists (a pure state as physicist say)], but was different in the past and also will be different in the future”[35]. It is clearly this camp that to be more friendly to the concept of creative time explained above.
@@
It seems that all A-theorists accept the theses of temporalism, i.e., that propositions change their truth-value over time, and of temporal disparity, i.e., the view that there is a metaphysical distinction between past, present and future [36]. The biggest disagreement is about which specific kind of temporal disparity one envisions. There are three main proposals: (i) Presentism [37, 38], which maintains that ``````` `no objects exist in time without being present” [35]; (ii) the Moving spotlight theory [39], an eternalist A-theory which takes events to be set on a fix manifold, but the present is singled out as if it was lit by a light moving along the manifold; and (iii) the Growing-block theory [40] (sometimes referred as the Growing-universe, or Growing-past), according to which events come into existence at present but, unlike in presentism, they persist in the past.
@@
The moving spot-light, as an eternalist theory, does not seem compatible with creative time, showing that there is not a one-to-one correspondence between A-theories and the concept of creative time. However, temporalism is an essential ingredient for creative time to process and certain A-theories are the natural metaphysical framework for describing the processing of creative time. Let us focus on the the Growing-block view. More precisely, the latter states that the present is metaphysically privileged because it is then that events become determinate (or ``````` `real” in the standard parlance of philosophy of time, which we do not endorse). But as new events become present, the past ones remain equally ``````` `real”, resulting in a growing block of reality [35]. In particular, we find parallels with a particular version of Growing-block theory, namely, E. Barnes and R. Cameron’s growing cloud of determinacy theory [41], which upholds that the openness of the future has nothing to do with its alleged non-existence (a common conclusion in philosphy of time). In the words of the authors, ``````` `the future is as yet unsettled. [T]hink of this ‘unsettledness’ with respect to future states of the world as a type indeterminacy. For all times $t_{1}$ and $t_{2}$ such that $t_{2}$ is later than $t_{1}$, it is indeterminate at $t_{1}$ what the state of the world is at $t_{2}$[41].
@@
This is the theory of the passage of time which seems most aligned with our concept of creative time. The future exists as a collection of real potentialities and the actualization of these potentialities is the act of creation that defines creative time. The present is therefore, for us, the edge of the events that turn from potential (the open future) to actual (the settled past). [^11] Although we have discussed the possibility that the remote past could again become indeterminate in Ref. [16]. So, while past, present and future are for us all in a sense real, the past is the (growing) collection of physical properties that got already actualized, the future is the collection of the potential properties, [^12] Note that some future properties have propensity one, hence are already actual [4]. and the present is the transition from potential to actual (see also [4]). [^13] Note that the present is not a global, universal property, but there are as many local presents as actualization events. Similar ideas were already put forward by H. Reichenbach who wrote: ``````` `The present, which separates the future from the past, is the moment when that which was undetermined becomes determined, and ’becoming’ means the same as ’becoming determined’.” [42]. Despite this, in Reichenbach, the present is still relative to someone or something, i.e. any moment can serve as the division between the determined past and the indeterminate future relative to who is simultaneous to that event (see [43] and reference therein for a discussion). Our proposal also has several similarities with S. McCall’s ``````` `objective time flow” [43]. Therein, the author considers that the universe has, at any point in time, a dynamical tree structure, with a single determined past and multiple possible futures. The concept of becoming is brought about by the update of the tree structure, when only one branch becomes determined. However, McCall’s proposal remains at an abstract level, whereas we present a naturalistic account of the indeterminate-actualization transition. Moreover, McCall does not conceive the possibility of objective (possibly biased) tendencies or propensities towards one possible future or another.
@@
To conclude this section, we find it worth positioning our contribution within the metaphysical debate on time. First of all, although we advocate some form of Growing-block theory, we notice that most existing accounts thereof in the philosophy of time have been developed with an a priori metaphysical approach. The novelty of our approach is to provide a naturalistic characterization of the passage of time and of the present, as an objective transition from the past to the future. Moreover, one of the main criticisms against the standard Growing-block theory (the so-called skeptic objection), which rendered this approach quite unpopular among philosophers, is that since both past and present have the same ontological status (i.e., they exist, as opposed to the future), how can we know that we live in the present and not in the past? It seems that our ideas expressed above overcome this skeptical objection exactly due to its naturalistic character. In fact, although both past and present are determinate (and therefore different from the future), they are also different from each other. The present is when creative time processes, namely, when the determination (or actualization) happens and new information gets created. The present has a dynamical nature, whereas the past remains fixed.
@@ VI.1 Relativity theory
As already recalled, geometric time is the conception of time derived from the regularities of some phenomena described by classical physics, like the motion of planets and satellites in the solar system which in turn led to a dominant deterministic world-view in physics. As remarked by G. E. M. Anscombe, ``````` `the high success of Newton’s astronomy was in one way an intellectual disaster: it produced an illusion… for this gave the impression that we had here an ideal of scientific explanation; whereas the truth was, it was mere obligingness on the part of the solar system, by having had so peaceful a history in recorded time, to provide such a model.” [44].
We agree with H. Reichenback that ``````` `the properties of time which the theory of relativity has discovered have nothing to do with its treatment as fourth dimension. This procedure was already possible in classical physics, where it was frequently used. However, according to the theory of relativity the four-dimensional manifold is of a new type; it obeys laws different from those of classical theory” [45]. And indeed it was the advent of the theory of (both special and general) relativity that popularized further the idea of geometric time. Therein, not only is time a parameter, but it seems to lose any special role with respect to space.
@@
Geometric Determinism Classical Non-tensed B-theory
Creative Indeterminism Intuitionistic Tensed A-theory
@@
Geometric time intervals are not relativistically invariant. Indeed, the (geometric) time of a moving observer at speed $v$ (with respect to a given reference frame) undergoes a dilation of a factor $\sqrt{1-\frac{v^{2}}{c^{2}}}$, according to the Lorentz transformations. Geometric time can be manipulated by using the scenario of the twin paradox. This means to prepare two identical systems ($K$ and $L$), send one of them (say $L$) away, and finally bring $L$ back together with $K$. Since $L$’s local clock passes at a slower rate than the one of $K$, the geometric time elapsed for the system $L$ will be shorter than the one elapsed for $K$ (the higher $v$, the larger the dilation). Let us consider again the example of the weather in Geneva in a month time, and let us assume that the same computation of the outcome is carried out with a computer $K$ and with a moving computer $L$ (assuming that they are using the same algorithm and identical computers). If at the event in which they meet again, $K$’s clock has registered a time $\Delta t^{\prime}$ (which is the time that it takes for her computer to return an outcome), $L$’s program would still be running due to time dilation. This shows again that geometric time can be manipulated. On the contrary, it is unclear whether there is any way to let creative time ``````` `run faster”.
@@
Yet, one can also think that the muon evolves deterministically while propagating through the atmosphere (thereby involving only geometric time) into a state of quantum superposition of decayed-undecayed and the process that actualizes the outcome ``````` `decayed” (which involves creative time) only happens inside the detector, hence at rest with Earth. In this way, creative time would not necessarily follow relativistic transformations. While this is admittedly a wild conjecture, it seems not to contradict the experimental evidence. This calls for further investigations.
@@
One might think that geometric time is more friendly to relativity than creative time (for a discussion on this see also [15]). Indeed, if there were only geometric time, then everything would be given all at once (determinism) and there would be no difference between time and space. But, actually, creative time might also be quite friendly to (special) relativity. Consider a true random number generator that produces (creates) a bit and let us fix the coordinate system such that this bit is produced at the origin. Along a time-like line passing through the origin, there is a time when the bit’s value is indeterminate and a time when it is determinate. Next, consider a space-like line in the future. Assuming that the bit has a determined value only inside the future light cone (see [15]), one can “move” along that space-like line and notice, similarly to moving along the time-like line, a segment thereof where the bit’s value is indeterminate and a segment where it is determinate. Note that along the time-like line there are only 2 regions –one of indeterminacy and one of determinacy–, while along the space-like line there are 3 regions, i.e., one can “move” from indeterminate to determinate and (without “returning”) again to indeterminate.
Finally, the idea that creative time fundamentally separates the past from the future may prima facie seem add odds with relativity. However, the events of actualization that characterize the development of creative time are to be considered local therefore there is no (global) simultaneity of the present, in accordance with relativity. The event of actualization from indeterminate to determinate exists only within the forward light-cone, while remaining indeterminate in any other region of spacetime (see [15] for a thorough discussion). Therefore distant regions have different ticking of creative time. However, one can conjecture that due to quantum nonlocality (of the propensities that characterize the indeterminacy), different ``````` `local presents” may be synchronised by the actualization at a distance of a nonlocal propensity (see also [4]): yet another wild conjecture.

### v2

@@ V Relation to the philosophy of time
In the vast literature devoted to the philosophy of time (see, e.g., [93, 74, 19, 53, 54] and references therein), one finds a main separation into two camps, following the seminal works of J. M. E. McTaggart [60]. The two camps are customarily referred to as A-theory and B-theory of time. While we acknowledge that the distinction in A- and B-theories presents a variety of modifications and it may not even encapsulate the most striking features of time in physics (e.g., directionality, openness of the future, objective becoming, etc.), we cannot avoid drawing some parallel with our distinction between creative and geometric time, respectively. [^19] Admittedly, there are a large variety of A- and B-theories. Here we believe we follow some widely accepted understandings thereof.
In fact, starting in reversed order, B-theorists regard positioning in time in dyadic relational terms: given two events $E_{1}$ and $E_{2}$, it is either the case that $E_{1}$ is before $E_{2}$, or simultaneous to $E_{2}$, or after $E_{2}$ [68]. Clearly, these relations are somehow ``static”, i.e., they are not tensed and hold unchanged since ever and forever.
@@
On the contrary, A-theorists uphold the view that characterizes the positioning of events in time with a monadic attribution: given an event $E$, it is either the case that $E$ is present, or past, or future [68].
Concerning the metaphysical status of time, the B-theory aligns more naturally with the spacetime structure of (deterministic) classical physics and relativity theory, where events are organized on a 4-dimensional manifold. Moreover, B-theorists tend to endorse Eternalism, i.e., the doctrine that all events are temporally on an equal footing and their relation of being in the past, present or future of each other merely reflects their placement on the manifold [40] (but, as we will see, this is not necessarily the case). Clearly, the concept of geometric time as expressed above fits best with the worldview offered by the B-theory camp.
A first point about Eternalism is that they seem to us to tacitly assume that events exist predeterminate, since all events necessarily are – in the static sense of being part of the block universe. As an example, consider two historically well-known events: the extinction of the dinosaurs and the French Revolution. In an indeterministic world, neither of these events was necessary; therefore, their before-after relation, as conceived in B-theories, would supposedly relate events that might never occur. At a remote past time, when both events – “dinosaurs’ extinction” and “French Revolution” – were merely potential, their before–after relation was likewise indeterminate, simply because the events themselves were not bound to happen. After the actualization of the dinosaurs’ extinction, but before the actualization of the French Revolution, one can rule out the possibility that the French Revolution is before (or simultaneous with) the extinction of the dinosaurs. However, it remains generally indeterminate whether the French Revolution will ever actualize (even if this indeterminacy may be characterized by a definite propensity [26]). [^20] We are not considering here the further complication that, from an evolutionary perspective, if the French Revolution occurs, this entails that humans have evolved, and thus the event must necessarily follow the extinction of the dinosaurs. Hence, B-theorists, who take the before–after relation as fixed, seem to us to assume that events are bound to occur.
Moreover, we note that the idea of ontic indeterminacy, which gives rise to an open future, is in tension with B-theories taken at face value. In fact, B-theorists deny the existence of the metaphysical properties of pastness, presentness, and futurity. For instance, in a greatly influential work to support a B-theory view in the philosophy of time, D. H. Mellor states that ``nothing really is past, present or future in itself”, and that ``tense is not an aspect of reality” [61]. In particular, we stress that B-theorists reject the idea that the future exists in itself. And by denying the existence of the future altogether, they also cannot accommodate the idea of an open future: There is no such a thing as the ``future” and, ipso facto, no ``open future”.
Hence, geometric time – which, as we have argued, is closely tied to determinism in physics – finds a more natural alignment with B-theory. In this view, there is no ontological distinction between past, present, and future, and nothing genuinely new ever comes into existence: no new information is added to the world.
A-theorists, on the contrary, accept the thesis of temporalism, i.e., that propositions change their truth-value over time, and of temporal disparity, i.e., the view that there is a metaphysical distinction between past, present and future [23]. This leads to the view that changes really happen: ``there is a way reality is (now, presently) which is complete [i.e., maximal given what exists (a pure state as physicist say)], but was different in the past and also will be different in the future”[40]. A-theory allows for a genuinely open future and fundamental indeterminism, which makes it the framework most naturally aligned with the notion of creative time.
@@
The biggest disagreement among different A-theories is about which specific kind of temporal disparity one envisions. There are three main proposals: (i) Presentism [57, 77, 94], which maintains that ``no objects exist in time without being present” [40]; (ii) the Moving spotlight theory [21], an eternalist A-theory which takes events to be set on a fix manifold, but the present is singled out as if it was lit by a light moving along the manifold; and (iii) the Growing-block theory [43] (sometimes referred as the Growing-universe, or Growing-past), according to which events come into existence at present but, unlike in presentism, they persist in the past.
@@
The moving spot-light, as an eternalist theory, does not seem compatible with creative time, showing that there is not a one-to-one correspondence between A-theories and the concept of creative time. The view we endorse here is akin in spirit to the Growing-block view, though with important distinctions. The growing block states that the present is metaphysically privileged because it is then that events become determinate (or ``real” in the standard parlance of philosophy of time, which we do not endorse). But as new events become present, the past ones remain equally ``real”, resulting in a growing block of reality [40]. In particular, we find inspiration in a variant of the Growing-block theory, namely, E. Barnes and R. Cameron’s growing cloud of determinacy theory [6], which upholds that the openness of the future has nothing to do with its alleged non-existence (a common conclusion in philosophy of time). In the words of the authors, ``the future is as yet unsettled. [T]hink of this ‘unsettledness’ with respect to future states of the world as a type of indeterminacy. For all times $t_{1}$ and $t_{2}$ such that $t_{2}$ is later than $t_{1}$, it is indeterminate at $t_{1}$ what the state of the world is at $t_{2}$[6]. As such, this theory aligns well with our physics-inspired idea that the future exists, metaphysically different from the past, as a collection of real potentialities and the actualization of these potentialities is the act of creation that defines creative time.
However, the Growing-block picture, along with the growing cloud of determinacy theory, is tenable only if we assume the existence of an absolute global present – a statement that holds in classical physics but is rejected by the theory of relativity. Therefore, our version of a growing-block universe incorporates the necessary modifications to remain compatible with relativistic principles. The present is defined, for us, as the edge of the events that turn from potential (the open future) to actual (the settled past). [^21] Although we have discussed the possibility that the remote past could again become indeterminate in Ref. [81]. For our view to remain compatible with relativity (see Section VI.1 and Ref. [25]), we must relax the absoluteness of determinate facts: the actualization of a propensity should itself be a relative property, defined with respect to regions of space–time, yet without invoking observers. Hence, a bit that becomes determinate does so only within its future light cone, while it remains indeterminate elsewhere. This results in multiple local presents, each defined by the collection of determination events lying on the edges of their past light cone. The set of local presents, however, contrarily to the Growing-block picture, cannot be collected into a global present.
We note that these considerations resonate with an approach to quantum gravity due to R. Sorkin, named causal sets [83, 32, 31]. This is an approach that takes spacetime as fundamentally made of discrete events related by partial order relations which encapsulate causality. The dynamics is a ``growth process” in which new elements come into existence from their causal ancestors.
Yet, our view allows to maintain that past, present and future are all in a sense real, metaphysically distinct, but relative to a light cone: the past is the (growing) collection of physical properties that got already actualized, the future is the collection of the potential properties, [^22] Note that some future properties have propensity one, hence are already actual [26]. and the present is the transition from potential to actual (see also [26]). Similar ideas were already put forward by H. Reichenbach who wrote: ``The present, which separates the future from the past, is the moment when that which was indeterminate becomes determinate, and ’becoming’ means the same as ’becoming determinate’.” [78]. In Reichenbach, however, the present is still relative to an observer or to a reference frame; that is, any moment can serve as the division between the determinate past and the indeterminate future, relative to whoever is simultaneous with that event (see [59] and references therein for a discussion).
@@
Our proposal also has several similarities with S. McCall’s ``objective time flow” [59]. Therein, the author considers that the universe has, at any point in time, a dynamical tree structure, with a single determinate past and multiple possible futures. The concept of becoming is brought about by the update of the tree structure, when only one branch becomes determinate. However, McCall’s proposal remains at an abstract level, whereas we present a naturalistic (physical) account of the indeterminate-actualization transition. Moreover, McCall does not conceive the possibility of objective (possibly biased) tendencies or propensities towards one possible future or another.
@@
We find it worth positioning our contribution within the metaphysical debate on time. First of all, the novelty of our approach is to provide a naturalistic characterization of the passage of time and of the present, as an objective transition from the past to the future. Moreover, one of the main criticisms against the Growing-block theory (the so-called skeptic objection), which rendered this approach quite unpopular among philosophers, is that since both past and present have the same ontological status (i.e., they exist, as opposed to the future), how can we know that we live in the present and not in the past? It seems that our ideas expressed above overcome this skeptical objection exactly due to its naturalistic character, based on indeterministic physics. In fact, although both past and present are determinate (and therefore different from the future), they are also different from each other. The present is when creative time processes, namely, when the determination (or actualization) happens and new information gets created. The present has a dynamical nature, whereas the past remains fixed.
To conclude, we have shown that geometric time is more naturally aligned with the worldview of B-theories, in which all events exist as well-determinate within a block universe. By contrast, creative time fits best with A-theories, which allow for an open future and thus fundamental indeterminism.
@@ VI.1 Relativity theory
As already recalled, geometric time is the conception of time derived from the regularities of some phenomena described by classical physics, like the motion of planets and satellites in the solar system which in turn led to a dominant deterministic world-view in physics. As remarked by G. E. M. Anscombe, ``the high success of Newton’s astronomy was in one way an intellectual disaster: it produced an illusion… for this gave the impression that we had here an ideal of scientific explanation; whereas the truth was, it was mere obligingness on the part of the solar system, by having had so peaceful a history in recorded time, to provide such a model.” [2].
We agree with H. Reichenback that ``the properties of time which the theory of relativity has discovered have nothing to do with its treatment as fourth dimension. This procedure was already possible in classical physics, where it was frequently used. However, according to the theory of relativity the four-dimensional manifold is of a new type; it obeys laws different from those of classical theory” [79]. And indeed it was the advent of the theory of (both special and general) relativity that popularized further the idea of geometric time. Therein, not only is time a parameter, but it seems to lose any special conceptual role with respect to space (besides the signature of the Minkowski metric, which assigns opposite signs to spatial and temporal components).
@@
Geometric Determinism Classical With LEM B-theory
Creative Indeterminism Intuitionistic Without LEM A-theory
@@
Geometric time intervals are not relativistically invariant. Indeed, the (geometric) time of a moving observer at speed $v$ (with respect to a given reference frame) undergoes a dilation of a factor $\sqrt{1-\frac{v^{2}}{c^{2}}}$, according to the Lorentz transformations. Geometric time can be manipulated by using the scenario of the twin paradox. This means to prepare two identical systems ($K$ and $L$), send one of them (say $L$) away, and finally bring $L$ back together with $K$. Since $L$’s local clock passes at a slower rate than the one of $K$, the geometric time elapsed for the system $L$ will be shorter than the one elapsed for $K$ (the higher $v$, the larger the dilation). Let us consider again the example of the weather in Geneva in a month’s time, and let us assume that the same computation of the outcome is carried out with a computer $K$ and with a moving computer $L$ (assuming that they are using the same algorithm and identical computers). If at the event in which they meet again, $K$’s clock has registered a time $\Delta t^{\prime}$ (which is the time that it takes for her computer to return an outcome), $L$’s program would still be running due to time dilation. This shows again that geometric time can be manipulated. On the contrary, it is unclear whether there is any way to let creative time ``run slower”.
@@
Yet, one can also think that the muon evolves deterministically while propagating through the atmosphere (thereby involving only geometric time) into a state of quantum superposition of decayed-undecayed and the process that actualizes the outcome ``decayed” (which involves creative time) only happens inside the detector, hence at rest with Earth, similarly to an atom in a superposition of an excited and a ground state. If the atom ``decays” (gets de-excited) then a photon is produced and the detection of the latter can demonstrate that the measurement took place. But if a photon is not detected, the atom can still be in superposition until it interacts with a detector. In this way, creative time would not necessarily follow relativistic transformations. While this is admittedly a wild conjecture, it seems not to contradict the experimental evidence. This calls for further investigations.
@@
One might think that geometric time is more friendly to relativity than creative time (for a discussion and proposed counterarguments see also [13, 25]). But, actually, creative time might also be quite friendly to (special) relativity. Consider a true random number generator that produces (creates) a bit and let us fix the coordinate system such that this bit is produced at a certain spacetime location, say the origin for simplicity. Along a time-like line passing through the origin, there is a time when the bit’s value is indeterminate and a time when it is determinate. Next, consider a space-like line that intersects the future light cone (see Fig. 2). Assuming that the bit has a determinate value only inside the future light cone (see [25]), one can “move” along that space-like line and notice, similarly to moving along the time-like line, a segment thereof where the bit’s value is indeterminate and a segment where it is determinate. Note that along the time-like line there are only 2 regions –one of indeterminacy and one of determinacy –, while along the space-like line there are 3 regions, i.e., one can “move” from indeterminate to determinate and (without “returning”) again to indeterminate.
Figure 2: Illustration of determinacy propagation within the light cone. Dotted lines indicate states of indeterminacy, while solid lines represent determinacy.
Finally, the idea that creative time fundamentally separates the past from the future may prima facie seem at odds with relativity. However, the events of actualization that characterize the occurrence of creative time are to be considered local therefore there is no (global) simultaneity of the present, in accordance with relativity. In Ref. [25], we have flashed out a proposal to make relativity and fundamental indeterminism compatible. This is to assume that event of actualization from indeterminate to determinate exists only within the forward light-cone, while remaining indeterminate in any other region of spacetime [25]. (In)determinacy is, therefore, itself relative, i.e., determinate facts are not absolute. This idea has gained considerable relevance in recent analyses in the foundations of quantum theory [14, 17, 28]. Our proposal, however, develops a theory of relative (determinate) facts with respect to regions of spacetime rather than to observers, as is the case in most of the above references. To understand this, consider for example two distant, space-like separated, observers, Alice and Bob, each with a local natural random number generator. If Alice’s random number generator actualizes a previously indeterminate bit to, say, the value $a=0$. For Bob, however, the value of the bit $a$ remains indeterminate until the intersection of the future light cones of Alice and Bob. In this light, the metaphysical property of being determinate is objective, but relative to a certain region of spacetime.
Therefore, different regions have different ticking of creative time. However, one can conjecture that due to quantum nonlocality (of the propensities that characterize the indeterminacy), different ``local presents” may be synchronized by the actualization at a distance of a nonlocal propensity (see also [26]). This admittedly remains a serious challenge and likely calls for a fully relativistic treatment of quantum nonlocality.